OpenStudy (anonymous):

If r and s are positive integers and 8^r=4^s, what is the value of s/r

OpenStudy (anonymous):

THIS IS SO HARD

OpenStudy (anonymous):

Have you studied logarithms?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay. Let's take a simpler approach. We know that \[8 = 2^3 ~~\rm{and} ~~ 4 = 2^2\] Therefore, we can rewrite the expression as\[2^{3r} = 2^{2s}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you follow how I went from \(8^r\) to \(2^{3r}\)?

OpenStudy (anonymous):

No c

OpenStudy (anonymous):

Okay, Recall that\[a^c \cdot a^d = a^{c \cdot d}\]

OpenStudy (anonymous):

oh yeah im starting to remember

OpenStudy (anonymous):

Sorry. That's wrong. \[a^c \cdot a^d = {a^{c + d}}\] We need \[(a^c)^d = a^{c \cdot d}\]

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

If we have \(8^r\) and we substitute \(2^3\) in for 8, we get \[\left ( 2^3 \right ) ^r\]This simplifies to \[2^{3r}\]

OpenStudy (anonymous):

Now, by Algebra magic, we can write \[2^{3r} = 2^{2s}\] as\[3r = 2s\] Can you find s/r now?

OpenStudy (anonymous):

what happens to the two big numbers

OpenStudy (anonymous):

I forget the proof, but let me try to explain. Let's say we have \[y^x = y^z\]Since y has the same value on both sides of the equal sign, we know that for \[y^x = y^z\] to be true, \[x = z\] has to be true. The same applies here, y = 2, x = 3r, and z = 2s 3r = 2s

OpenStudy (anonymous):

after 3r=2s what happens next

OpenStudy (anonymous):

Set up the ratio of s/r\[{s \over r} = {3 \over 2}\]If we cross-multiply, we end up back at 3r = 2s

OpenStudy (anonymous):

i see.and when we do cross-multiply what happens next. is 3r=2s the answer

OpenStudy (anonymous):

No. We want the ratio s/r. I was just noting how to check to make sure we had the right ratio. If we cross multiply and get the original expression, we are right. The answer is s/r = 3/2

OpenStudy (anonymous):

ohh okay thank you so much